Suppose a firm has its TR and TC functions estimated as follows:



TR = 300Q-3Q2



TC = 500 + 50Q + 2Q2



Find (a) profit function of the firm, (b) the quantity of output (Q) that maximizes the firm's profit.




1
Expert's answer
2022-01-17T16:23:26-0500

a)

"PF=TR-TC=-500+250Q-5Q^2,"

b)

"PF'=250-10Q,"

"Q=25,"

"PF=2625."